If the 7th term of g.p. be 16 then the product of first 13 terms is?

Grade:12

1 Answers

Meet
6 years ago
a(r)^6=16 then product of first 13 terms is a.(ar).(ar^2)......(ar^12) = a^13.(r^6)^13 and its equal to [a(r)^6]^13 = 16^13